OpenStudy (anonymous):

f(x)=4x-7 a.f(3) b.f(-5)

TheSmartOne (thesmartone):

What are we trying to do with it? Graph it?

OpenStudy (anonymous):

no solve it with this (a).f(3) b.f(-5)

TheSmartOne (thesmartone):

Ok a. We have to plug in 3 for x So it becomes f(3)=4(3)-7 f(3)=12-7 f(3)=5

TheSmartOne (thesmartone):

Can you solve the second one and I will see if you understood it.

OpenStudy (anonymous):

-27

TheSmartOne (thesmartone):

Tell me what you got and then if you got it wrong I will explain it to you. Open Study is not for just getting the answers without understanding it.

TheSmartOne (thesmartone):

Exactly -27 is correct

OpenStudy (anonymous):

are they both right .i though there coudnt be 2 answers